OpenStudy (anonymous):

[Sqrt(32cos^2x+1)-4]/[4-sqrt(32sin^2x+1)] Solve.

OpenStudy (anonymous):

Well for solving first we need an equation ...

OpenStudy (anonymous):

\[(\sqrt{32\cos^2x+1}-4)/(4-\sqrt{32\sin^2x+1})\]\[(\sqrt{32\cos^2x+1}-4)/(4-\sqrt{32\sin^2x+1})=1\]

OpenStudy (anonymous):

Sorry, my bad.

OpenStudy (experimentx):

let sin^2 x = u, we will have quadratic equation in u